OpenStudy (anonymous):

Use the formula for future value with simple interest to find the missing quantity. Round to the nearest cent. A = $16,000, P = ?, r = 6.5%, t = 3 years I got 780.49. Is this correct?

OpenStudy (mr.math):

Try again!

OpenStudy (anonymous):

Oh man.

OpenStudy (mr.math):

Could you write down the formula you used and how you used it.

OpenStudy (mr.math):

?

OpenStudy (anonymous):

I used A = P(1+rt) 16000= p (1.6.5 *3)

OpenStudy (mr.math):

You mean \(16000=P(1+0.065\times3)\).

OpenStudy (mr.math):

Be careful when dealing with the interest rate. 6.5%=0.065.

OpenStudy (mr.math):

3%=3/100=0.03. 4%=4/100=0.04 5.5%=5.5/100=0.055 and so on.

OpenStudy (anonymous):

ohhh... okay THANK YOU! so for this question: Determine the present value P you must invest to have the future value A at simple interest rate r after time t. Round to the nearest cent. A = $7,500, r = 9.5%, t = 2 years Would I do 7,500 = P (1 + 0.095 x 2) and then the answer would be 6,302.50?

OpenStudy (mr.math):

Great!

OpenStudy (espex):

Correct.

OpenStudy (anonymous):

THANK YOU!